# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ D.2 Local Sublinear Convergence
Theorem D.7 (Local linear convergence by the away-step FW algorithm).
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $\omega:=\frac{(\mu-\rho)^{2}}{8\mu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\frac{(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}}{2LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma D.6, we have
$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (D.8)
where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. From Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
@@
where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma 3.3 and the quadratic growth condition of $f$) and (D.8).
In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$. For away steps, it seems that we only obtain a monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from (D.5) and $\rho<L$, we obtain
@@
In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest. The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from (D.5) and $\rho<L$. We use Lemma B.11 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim. ∎
@@
D.5 Local Linear Convergence over Uniformly Convex Sets
In the convex optimization case, Canon and Cullum [Canon1968-sr] established an early lower bound on the convergence rate of the FW algorithm. However, in the special case of $P$ being strongly convex, Garber and Hazan [Garber2015-de] showed that one can improve upon that lower bound. Kerdreux et al. [Kerdreux2021-kc] establish it in the case of $P$ being uniformly convex. We will now carry over this result to establish local linear convergence for the case where $f$ is weakly convex and $L$-smad on $P$.
@@ D.5 Local Linear Convergence over Uniformly Convex Sets
Definition D.8 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
[…]
@@
Proposition D.9 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
@@
Theorem D.10 (Local linear convergence over uniformly convex sets).
@@
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma 3.4 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition D.9, we have
@@
where the third inequality holds from Lemma 3.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (D.5). For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
@@
Thus, we have the claim. For $p>2$, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (D.5). In that case, a local linear rate in Theorem D.10 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem D.11 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
@@
Using an argument similar to Theorem D.10, we obtain
@@
where the third inequality holds from the PL inequality (B.6) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎
Appendix E Experiments
@@
  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy [Pedregosa2020-ay]
@@
  • MD: the mirror descent [Nemirovskij1983-sm]
E.1 Nonnegative Linear Inverse Problems
Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$. This class of problems has been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de] as well as optimization [Bauschke2017-hg, Takahashi2024-wz]. Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate. From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$. Recovering $x$ can be formulated as a minimization problem:
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (E.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence. Problem (E.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm]. The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$ (random seed 1234). All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 5 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
Figure 5: Nonnegative linear inverse problem for $(m,n)=(100,1000)$.
Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$. The results are reported as mean $\pm$ standard deviation.

### v4

@@ E.2 Local Sublinear Convergence
Proof of Theorem 5.4.
By letting $h_{t}=f(x_{t})-f^{*}$ and using Lemma E.6, we have
$h_{t}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x_{t}),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}\leq\frac{8\mu\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{(\mu-\rho)^{2}\delta^{2}}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle^{2}}{\omega\delta^{2}},$ (E.7)
where the second inequality holds because $d_{t}$ is either $x_{t}-v_{t}^{\textup{FW}}$ or $v_{t}^{\textup{A}}-x_{t}$ with $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle/2$ in Lines 3 and 3 of Algorithm 3. From Lemma B.8 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), we obtain
@@
where the second inequality holds because of $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}-\frac{\rho}{2}\|x_{t}-v_{t}\|^{2}\geq\left(1-\frac{\rho}{\mu}\right)h_{t}$ (from Lemma C.2 and the quadratic growth condition of $f$) and Equation E.7.
In $\nu=1$, for Frank–Wolfe steps, we have $\gamma_{t,\max}\left(1-\frac{\rho}{\mu}\right)=\left(1-\frac{\rho}{\mu}\right)\geq\frac{1}{8}\left(1-\frac{\rho}{\mu}\right)^{2}\frac{\delta^{2}}{D^{2}}=\frac{\omega\delta^{2}}{\mu D^{2}}\geq\frac{\omega\delta^{2}}{2LD^{2}}$ by $0<\rho<\mu\leq L$ and $\delta\leq D$. For away steps, it seems that we only obtain a monotone progress $h_{t+1}<h_{t}$ because it is difficult to estimate $\gamma_{t,\max}$ below. However, $\gamma_{t,\max}$ cannot be small too often. Let us consider $\gamma_{t}=\gamma_{t,\max}$ in an away step. In that case, $v^{\textup{A}}_{t}$ is removed from the active set $\mathcal{S}_{t+1}$ in line 3. Moreover, the active set can only grow in Frank–Wolfe steps (line 3). It is impossible to remove more vertices from $\mathcal{S}_{t+1}$ than have been added in Frank–Wolfe steps. Therefore, at most half of iterations until $t$ iterations are in away steps, i.e., in all other steps, we have $\gamma_{t}=\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{2LD_{\phi}(v_{t},x_{t})}<\gamma_{t,\max}$ and then $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$. Because we have $h_{t+1}\leq\left(1-\frac{\omega\delta^{2}}{4LD^{2}}\right)h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest, using $h_{1}\leq 2LD^{2}$ from Equation E.5 and $\rho<L$, we obtain
@@
In $\nu\in(0,1)$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right),\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{1-\nu}{2\nu}}\right\}\cdot h_{t}$ for at least half of the iterations and $h_{t+1}\leq h_{t}$ for the rest. The initial bound $f(x_{1})-f^{*}\leq(\rho+L)D^{2}\leq 2LD^{2}$ holds from Equation E.5 and $\rho<L$. We use Lemma C.11 with $c_{0}=2LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{(\omega\delta^{2})^{\frac{1+\nu}{2\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}$ and obtain the claim. ∎
@@
E.5 Local Linear Convergence over Uniformly Convex Sets
In the convex optimization case, Canon and Cullum (1968) established an early lower bound on the convergence rate of the FW algorithm. However, in the special case of $P$ being strongly convex, Garber and Hazan (2015) showed that one can improve upon that lower bound. Kerdreux et al. (2021b) establish it in the case of $P$ being uniformly convex. We will now carry over this result to establish local linear convergence for the case where $f$ is weakly convex and $L$-smad on $P$.
@@ E.5 Local Linear Convergence over Uniformly Convex Sets
Definition E.7 ($(\alpha,p)$-uniformly convex set (Braun et al., 2025, Definition 2.18)(Kerdreux et al., 2021b, Definition 1.1)).
Let $\alpha$ and $p$ be positive numbers. The set $P\subset\mathbb{R}^{n}$ is $(\alpha,p)$-uniformly convex with respect to the norm $\|\cdot\|$ if for any $x,y\in P$, any $\gamma\in[0,1]$, and any $z\in\mathbb{R}^{n}$ with $\|z\|\leq 1$ the following holds:
@@
Proposition E.8 (Scaling inequality (Braun et al., 2025, Proposition 2.19)(Kerdreux et al., 2021b, Lemma 2.1)).
@@
Theorem E.9 (Local linear convergence over uniformly convex sets).
@@
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma B.8 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition E.8, we have
@@
where the third inequality holds from Lemma C.2 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and Equation E.5. For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
@@
Thus, we have the claim. For $p>2$, we use Lemma C.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from Equation E.5. In that case, a local linear rate in Theorem E.9 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem E.10 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
@@
Using an argument similar to Theorem E.9, we obtain
@@
where the third inequality holds from the PL inequality Equation C.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma C.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎
Appendix F Experiments
@@
  • EucFW: the FW algorithm with the adaptive (Euclidean) step-size strategy (Pedregosa et al., 2020)
  • EucAFW: the away-step FW algorithm with the adaptive (Euclidean) step-size strategy (Pedregosa et al., 2020)
@@
  • MD: the mirror descent (Nemirovskij and Yudin, 1983)
  • ProjGD: the projected gradient descent algorithm (see, e.g., Beck (2017))
F.1 Nonnegative Linear Inverse Problems
Given a nonnegative matrix $A\in\mathbb{R}_{+}^{m\times n}$ and a nonnegative vector $b\in\mathbb{R}_{+}^{m}$, the goal of nonnegative (Poisson) linear inverse problems is to recover a signal $x\in\mathbb{R}_{+}^{n}$ such that $Ax\simeq b$. This class of problems has been studied in image deblurring (Bertero et al., 2009) and positron emission tomography (Vardi et al., 1985) as well as optimization (Bauschke et al., 2017; Takahashi and Takeda, 2025). Since the dimension of $x$ is often larger than the number of observations $m$, the system is indeterminate. From this point of view, we consider the constraint $\Delta_{n}:=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}\leq 1\}$. Recovering $x$ can be formulated as a minimization problem:
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (F.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the KL divergence. Problem Equation F.1 is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from (Bauschke et al., 2017, Lemma 8).
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD(Nemirovskij and Yudin, 1983). The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by (Beck, 2017, Example 3.71), and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$ (random seed 1234). All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 5 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows the average values of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
Figure 5: Log-log plot of primal and FW gaps on nonnegative linear inverse problem for $(m,n)=(100,1000)$.
Table 2: Average values of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$. The results are reported as mean $\pm$ standard deviation.